What is the density of Carbon Dioxide gas at Standard Temperature and Pressure? 20 degree Celsius and 735 millimeter Mercury?


Note: density = mass/Volume       n = mass/molar mass  


1
Expert's answer
2021-01-27T04:39:14-0500

V= nRT/p= 1 x 62400 x 293/735 = 24 875,102 ml.=24,875102 L.

density=44/24,875102 = 1,76883697 g/L.